find the area of the region between the graph of y = x^3 and the x-axis from x= 1 to x= 4
Answer by Alan3354(69443)

find the area of the region between the graph of y = x^3 and the x-axis from x= 1 to x= 4
------------
The integral of x^3 = x^4/4 + C
At x = 4: x^4/4 = 64
At x = 1: x^4/4 = 1/4
Area = 64 - 1/4 = 63.75 sq units